# RoboGrind: Intuitive and Interactive Surface Treatment with Industrial Robots

### 26 Feb 2024 v1

@@ III-3 Force-Controlled Surface Treatment
RoboGrind performs program generation and path planning based on a 3D point cloud representation of the surface. We use a Gocator 2490111LMI Technologies GmbH, Teltow, Germany laser line scanner with a field of view of up to 2 m and a resolution of 6 $\mu$m. The starting point of a scan is defined manually according to the measurement range. Segmentation, outlier detection, and downsampling are performed. Based on an initial calibration step, the point cloud is transformed into the robot’s base frame. We propose an automatic defect detection pipeline that combines two outlier detection algorithms: First, curve fitting by linear regression is performed separately for the points of each scanned line and, based on several thresholds, outliers and defects are marked and excluded. Second, a StatisticalOutlierRemoval filter from the Point Cloud Library (PCL) based on k-Nearest Neighbors (k-NN) is used to detect points with few neighbors. This allows detecting scanning artefacts that cause holes in point clouds, and eliminates outliers that do not belong to a scanned part. It also helps reducing the number of false positives detected, as these have no high point density and have too few neighbors. The combined algorithms can detect all defects that deviate from an ideal surface shape, including dents, bumps, scratches, and rough surface areas. Unlike neural networks, they require no training.
@@ VIII-A1 Perception
Evaluation of laboratory experiments regarding the perception and damage detection demonstrated that perception is robust in over 71 % of the 24 samples and 96 included damages. 69 defects were correctly detected, 27 were missed and 7 marked spots were false positives. Remaining problems are holes in point clouds that are caused by failed measurements and false positive defect detections. Failed measurements result from darker colored or dirty surfaces that are not detectable by the used laser line scanner. False positives occur in sections where the depth of the defects is relatively low because of the dynamic thresholds of the algorithms.
@@ Acknowledgment
This work was partially supported by the German state of Baden-Württemberg in the project RoboGrind under the grant no. BW1_0079/01.

### 27 Feb 2024 v2

@@ III-3 Force-Controlled Surface Treatment
RoboGrind performs program generation and path planning based on a 3D point cloud representation of the surface. We use a Gocator 2490111LMI Technologies GmbH, Teltow, Germany laser line scanner with a field of view of up to 2 m and a resolution of 6 $\mu$m. The starting point of a scan is defined manually according to the measurement range. Segmentation, outlier detection, and downsampling are performed. Based on an initial calibration step, the point cloud is transformed into the robot’s base frame. We propose an automatic defect detection pipeline that combines two outlier detection algorithms: First, curve fitting by linear regression is performed separately for the points of each scanned line and, based on several thresholds, outliers and defects are marked and excluded. Second, a StatisticalOutlierRemoval filter from the Point Cloud Library (PCL) based on k-Nearest Neighbors (k-NN) is used to detect points with few neighbors. This allows detecting scanning artefacts that cause holes in point clouds, and eliminates outliers that do not belong to a scanned part. It also helps reducing the number of false positives detected, as these have no high point density and have too few neighbors. The combined algorithms can detect all defects that deviate from an ideal surface shape, including dents, bumps, scratches, and rough surface areas. Unlike neural networks, they require no training. The code can be found in our GitLab repository [^2] https://gitlab.com/rahm-lab/robogrind-surfacetreatment.
@@ VIII-A1 Perception
Evaluation of laboratory experiments regarding the perception and damage detection demonstrated that perception is robust in over 71 % of the 24 samples and 96 included damages. 69 defects were correctly detected, 27 were missed and 7 marked spots were false positives. Remaining problems are holes in point clouds caused by failed measurements and false positive defect detections. Failed measurements result from darker colored or dirty surfaces that are not detectable by the laser scanner. False positives occur in sections where the depth of the defects is low because of the dynamic thresholds of the algorithms.
@@ Acknowledgment
This work was supported by the state of Baden-Württemberg in the project RoboGrind under grant BW1_0079/01, the DFG CRC EASE (CRC #1320) and the EU project euROBIN (grant 101070596).